Question: out of 900 students at Ross middle school, 33% are in choir. How many students are in choir?
I know that if 33 is over 100, I need to find 'x' over 900 but I don't know how!

Answer by rfer(16322)

just multiply
900*.33=297
----------------
or cross multiply, then divide
x/900=33/100
100x=29700
29700/100=297